OpenStudy (anonymous):

A girl went to a horse race and she forgot to count the horses and humans there, but she knows there are 196 legs and 74 heads. How many horses and humans were there at the horse race?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

YOU ARE WRONG

OpenStudy (anonymous):

Let x be the number of horses and y be the number of humans. \[x+y=74\] \[4x+2y=196\] Use substitution method. \[x=74-y\] \[4(74-y)+2y=196\] \[296-4y+2y=196\] \[2y=100\] \[y=50\] Therefore the number of humans present was 50. To find the number of horses (x-value), substitute the y-value into the equation x=74-y

OpenStudy (anonymous):

25 horses?

OpenStudy (anonymous):

Man today's not your day son. Have you learnt how to subtract because that's a pretty easy subtraction there and you mucked it up.

OpenStudy (anonymous):